A and B are Boolean random variables. Given: P(A=True) = 0.3, P(A=False) = 0.7, P(B=True|A=True) = 0.4, P(B=False|A=True) = 0.6, P(B=True|A=False) = 0.6, P(B=False|A=False) = 0.4 . Calculate P(A=True|B=False) by Bayes rule.

Grade:12th pass

2 Answers

abhay
6 years ago
answer is 0.39, using bayes rule NPTEL assign 4 right?, P(A=True|B=False)=P(B=False|A=True)*P(A=True)/P(B=False|A=True)*P(A=True)+P(B=True|A=True)*P(A=False)
Shubham Sharma
2 years ago
A and B are Boolean random variables. Given: P(A=True)=0.3,P(A=False)=0.7,P(B=True|A=True)=0.4,P(B=False|A=True)=0.6, P(B=True|A=False)=0.6, P(B=False|A=False)=0.4. Calculate P(A=True|B=False) by Bayes rule.